OpenStudy (anonymous):

Determine the average rate of change of the function, f(x)=3^x over the interval x=3 to X=4.

OpenStudy (anonymous):

my guess here is from 3 to 4, you take the slope formula because a slope is a rate of change from one number to another, so take (y2-y1)/(x2-x1)= m (3^4-3^3)/(4-3) = 81-27/1=54

OpenStudy (anonymous):

where did you get 3^3 and 3^4

hero (hero):

@Sukhvir, all you have to do is find both (x,y) coordinates. Remember y = f(x) and if you have any two points, you can calculate the slope using the slope formula.

OpenStudy (anonymous):

but i only have two x's?

OpenStudy (anonymous):

y1-y2/3-4 ?

hero (hero):

Remember, y = f(x) therefore y = 3^x input the x's into that formula to find the y's. No more clues for you.

OpenStudy (anonymous):

y=3^(3-4)

OpenStudy (anonymous):

y=3^(1)

OpenStudy (anonymous):

therefor y=3

hero (hero):

Why do you keep doing exactly the wrong thing?

hero (hero):

Input the x's into the formula ONE AT A TIME. You have TWO x's. You need TWO y's

OpenStudy (anonymous):

so i am using y2-y1/x2-x1

OpenStudy (anonymous):

or y=mx+b

hero (hero):

Input the x's into the formula y = 3^x ONE AT A TIME to get: y = 3^(3) y = 3^(4) I can't believe I had to make it that obvious.

hero (hero):

Let's see if you can actually calculate the y values

OpenStudy (anonymous):

27-81/3-4

OpenStudy (anonymous):

=54

hero (hero):

Would it not be easier to do \(\frac{81 - 27}{4-3}\)

OpenStudy (anonymous):

y=(-54)

hero (hero):

Anytime, you can avoid dealing with negatives, you should avoid.

hero (hero):

81 - 27 = 54 4 - 3 = 1 54/1 = 54 You had it right the first time. I just meant for you to avoid the negatives when necessary.

OpenStudy (anonymous):

ok thanks a lot man really appreciated

hero (hero):

Why do you confuse yourself so much?

hero (hero):

Confusing yourself is not necessary to solve math problems.

OpenStudy (anonymous):

im not that confident with my math skills i

OpenStudy (anonymous):

i question my methods

hero (hero):

You should.